OpenStudy (anonymous):

Which of the following functions satisfies the following conditions: f is increasing and has an x intercept at x=0? a. (e^x)+2 b.2-(2e^x) c.2-(e^-x) d.(-e^-x) e.(e^2x)-1

OpenStudy (anonymous):

It's option e. (e^2x) is an increasing function as it has a positive slope. And to get the x-intercept, put y = 0 and solve for x. It will be: y = (e^2x) - 1 0 = (e^2x) - 1 e^2x = 1 2x = ln (1) x = 0 (because ln (1) = 0) so that means the x-intercept is 0.